1/2

Rabindranath Tagor,  Rabindranath Tagor,
The young prince  Shubhokkan

tradukita de Rabindranath Tagor

O mother, the young Prince is to pass by our

door, – how can I attend to my work this morning? Show me how to braid up my hair; tell me what garment to put on. Why do you look at me amazed, mother? I know well he will not glance up once at my

window; I know he will pass out of my sight in the

twinkling of an eye; only the vanishing strain of the flute will come sobbing to me from afar. But the young Prince will pass by our door, and I will put on my best for the moment.

Ogo maa, Raajaar dulaal jaabe aaji mor gharer shomukhopathe, aaji e probhaate grihokaaj loye rohibo balo ki mate. Bole de aamaae ki koribo shaaj, ki chaaxde kabori bexdhe labo aaj, poribo angge keamon bhangge kon baroner

baash. Maa go, ki holo tomaar, abaak nayone mukho-

paane keano caash. Aami daaxrxaabo jethaae baataayonkone she caabe naa shethaa jaani taahaa mone – phelite nimesh deakhaa habe shesh, jaabe she

shudur pure, shudhu shangger baaxshi kon maatxh hote baa-

jibe beakul shure. Tobu raajaar dulaal jaabe aaji mor gharer sho-

mukhopathe, shudhu nimesh laagi naa koriyaa besh rohibo balo

ki mate.

O mother, the young Prince did pass by our

door, and the morning sun flashed from his chariot. I swept aside the veil from my face, I tore the ruby chain from my neck and flung it in his path. Why do you look at me amazed, mother? I know well he did not pick up my chain; I know it was crushed under his wheels leaving a red stain upon the dust, and no one knows what my gift was nor to

whom. But the young Prince did pass by our door, and I flung the jewel from my breast before his

path.

Ogo maa, raajaar dulaal gealo coli mor gharer

homukhopathe, probhaater aalo jholilo taahaar sharnoshikhar

rathe. Ghomtxaa khashaaye baataayon theke nimesher laagi niyechi maa dekhe, chixrxi monihaar phelechi taahaar pather dhu-

laar ’pare. Maa go, ki holo tomaar, abaak nayone caahish

kisher tare! Mor haar-chexrxaa moni neay ni kurxaaye, rather caakaae geache she guxrxaaye caakaar cinho gharer shomukhe porxe aache

shudhu aaxkaa. Aami ki dilem kaare jaane naa she keu – dhulaae

rohilo dxhaakaa. Tobu raajaar dulaal gealo coli mor gharer

shomukhopathe– mor bokkher moni naa pheliyaa diyaa rohibo balo

ki mate.

…  …

2/2

Traduko de la bengala poemo “Shubhokkan” de Rabindranath Tagor (∗1861-05-07 – †1941-08-07) en la Anglan de Rabindranath Tagor (∗1861-05-07 – †1941-08-07).

Arg-844-1700 (2013-02-26 23:27:58)

Tiu estas angligo, kiun la poeto verkis mem laŭ sia origine en la bengala lingvo verkita teksto. (Konfirmis tion al mi, Manfred Retzlaff, s-ro Probal Dasgupta.) Vidu ankaŭ la retejojn http: // eo. wikipedia. org/ wiki/ Probal_Dasgupta kaj http: // www. recmusic. org/ lieder/ get_text. html? TextId= 18070 .

Verkinto de tiu ĉi bengala poemo estas Rabindranath Tagor (∗1861-05-07 – †1941-08-07).

Arg-844-1699 (2013-02-26 23:16:53)

Bengala-lingva teksto (en latina-skriba transskribo) verkita de la poeto Rabindranath Tagor / Robindronath Tagor, (07.05.1861 - 07.08.1941). Tiun ĉi tekston mi, Manfred Retzlaff, ricevis de s-ro Probal Dasgupta per ret-mesaĝo en 2011-02-14. Ĝi estas prenita el la plena verkaro de la poeto Robindronath Tagor eldonita de Visvabharati. Vidu ankaŭ la Esperantigon faritan de Kálmán Kalocsay, kiu aperis en 1950 en la Esperanta revuo “La Praktiko”, kaj kiun enigis la dana samideano Lars Kromann en sian retan kolekton http: // www. literaturo. dk sub http: // www. literaturo. dk/ junaprinco. htm . .

http://www.poezio.net/version?poem-id=844&version-id=1700,1699